# Photodynamic melting of phase-reversed charge stripes and enhanced condensation

### v1

@@ Abstract
The interplay between charge stripes and pairing is a longstanding point of scrutiny in a broad class of unconventional superconductors since, in some cases, it is unclear whether their intertwining benefits the ensuing superfluidity. Experiments that explore the out-of-equilibrium dynamics of these systems try to tip the balance in favor of one phase or the other by selective coupling to relevant modes. Leveraging the fact that competition between stripes and pairing is not exclusive to fermionic systems, we explore the photoirradiation dynamics of interacting hardcore bosons, in which density wave phase-reversal melting gives rise to enhanced superfluid properties, quantified by the dynamic amplification of zero-momentum occupancy and charge stiffness. Our results, obtained using unbiased methods for an interacting system on a ladder geometry, demonstrate how one can engineer time-dependent perturbations to release suppressed orders, potentially providing insight into the underlying mechanism in related experiments.
Introduction.—  Time-dependent experiments provide another angle in the longstanding quest to understand the mechanism binding otherwise repulsive fermions, supporting the emergence of unconventional superconductivity. In particular, rapid energy insertion in the form of an ultrafast laser pump has arisen as a capable technique [1, 2] that coherently drives the dynamics of the electrons in ways that can control magnetic [3, 4], charge [5, 6, 7, 8], and superconducting orders [9, 10, 11, 12, 13, 14, 15]. Owing to the entwining of the different orders that permeate the doped cuprate physics [16, 17], such an approach is one in which coupling to certain degrees of freedom can suppress specific ordering while releasing others [2, 18].
@@
Although the physics of stripes in equilibrium, including fluctuating/incommensurate ones, is now well studied within a broad class of models and numerical methods [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], direct indication of the dynamical interplay between charge ordering and pairing is nonexistent in unbiased numerics of many-body models even though this is the precise mechanism explored in experiments [9, 43, 11, 44] – we aim to fill this void  [^1] We note microscopic models for cuprates have been studied under photoirradiation [84, 85, 86, 87], but the interplay with stripe physics is currently missing.. Our key observation is that the stripe physics and antiphase of local order are not intrinsic to repulsive interactions [46], nor to fermionic statistics [47], including the fluctuating-static dichotomy [47]. This fundamental, and highly non-trivial, analogy allows one to unbiasedly solve the dynamics of Hamiltonians that contain specific ingredients of actual cuprate physics — namely, strong interactions, superfluidity, and antiphase correlations across stripes — even if the pairing is isotropic, unlike what is observed in a large class of unconventional superconductors.
Our main result is that by adjusting the characteristics of the time-dependent perturbation in the form of a pump in a strongly correlated system, one can show that an effective nonlinear optical coupling is responsible for exciting the system from the ground state in a way that a static stripe with antiphase ordering melts while (quasi-)superfluidity is enhanced dynamically.
Model.—  Our starting point is the Hubbard model for repulsive hardcore bosons,
@@
Figure 1: (a) Schematic illustration of Eq. (1) with relevant terms annotated, subjected to $x$-polarized photoirradiation. Here, the hardcore bosons are represented by a composite local fermion. Momentum distribution profiles $n_{\bf k}$ focusing on $k_{y}=0$ (b) and $k_{y}=\pi$ (c), comparing the equilibrium (before pump) with the out-of-equilibrium (after pump, averaged results from $t\in[11t_{d},26t_{d}]=[72,169]t_{h}^{-1}$; shading gives the standard deviation) response to the external field. We observe an average relative enhancement of about 37% for the zero-momentum condensation under the photoirradiation. (d) The $x$-direction charge stiffness resolved for the first 200 eigenstates $|m\rangle$ of $\hat{\cal H}$; marker colors are mapped by the overlap of the time evolved state at $t_{1}=5t_{d}$ and the corresponding $|m\rangle$. (e) The instantaneous energy $E(t)=\langle\psi(t)|\hat{\cal H}(t)|\psi(t)\rangle$ over the dynamics; the shaded region marks the probing for the dynamical charge stiffness starting at $t_{1}$, with a linear in time vector potential with slope $\delta A_{x}$ – top panel depicts the total ${\bf A}(t)$. (f) Energy variation in the probe regime as a function of the effective flux $\Phi_{x}(t)=\delta A_{x}L_{x}(t-t_{1})$ for $t>t_{1}$$\delta A_{x}=0.01$. Dashed line displays an approximant at small fluxes that quantifies the dynamical charge stiffness at $t=t_{1}$, confirming a robust charge transport induced by the pump (see text). Pulse parameters are $\Omega=1.713\ t_{h}$, $A_{0}=0.62$ and $t_{d}=6.5t_{h}^{-1}$.
@@
Enhanced condensation and dissipationless transport.—  Interacting hardcore bosons at $T=0$ give rise to either charge ordered phases or a superfluid phase [52, 53, 54, 55, 56, 57], depending on the interaction-hopping ratio $V/t_{h}$ — in what follows we set $V/t_{h}=4$. Tendency towards phase separation and stripe formation in the vicinity of half-filling, $\langle\hat{n}\rangle=1/2$, has been reported [55, 57], and stabilization of such a regime can be made explicit via an external periodic potential [47] in the doped regime  [^4] Including an extra potential has been used in other contexts to stabilize magnetic ordering for example [88]. Notably, in this scenario, the propensity for spatial organization of holes is accompanied by a reversal in the phase of the checkerboard density modulation across a stripe boundary provided the corresponding chemical potential $V_{0}$ is sufficiently large – this is often dubbed as a $\pi$-phase shift [33, 46, 47]. Focusing on the hole density $\delta\equiv 1/2-\langle\hat{n}\rangle=1/16$, Fig. 2(a) displays the connected density-density correlations across a stripe, $c_{{\bf i}_{0}}(dx,dy)=\langle\hat{n}_{{\bf i}_{0}}\hat{n}_{{\bf i}+d{\bf r}}\rangle-\langle\hat{n}_{{\bf i}_{0}}\rangle\langle\hat{n}_{{\bf i}+d{\bf r}}\rangle$; ${\bf i}_{0}$ is a reference site and $d{\bf r}\equiv(dx,dy)$ the spatial displacement. As $V_{0}/t_{h}$ increases, the same sublattice correlation $c_{{\bf i}_{0}}(2,0)$ across a stripe transitions from positive to negative row whereas $c_{{\bf i}_{0}}(2,1)$ shows the opposite trend, establishing a direct verification of the antiphase charge ordered pattern once hole stripes are sufficiently robust.
@@
Nonetheless, condensation need not be necessarily tied to superfluid behavior, in general [60]. A necessary condition for dissipationless charge transport  [^6] But not sufficient for a finite superfluidity, which entails phase coherence, quantified by a finite superfluid weight. is given by the charge stiffness (or Drude weight), whose finite value indicates an ideal metal with ballistic transport [62, 63, 64]. For the $m$-th eigenstate of $\hat{\cal H}$ with eigenenergy $E_{m}$,
@@
Figure 1(f) shows the time-dependence of the energy gain during the probe time. Indeed, at short times, a quadratic dependence in $\Phi_{x}(\Delta t)$ emerges, indicating a finite $D^{(x)}$ and thus dissipationless transport after the pulse. The second estimation of this physical behavior is that $D^{(x)}$ closely matches with an approximant of the out-of-equilibrium Drude weight $\tilde{D}^{(x)}=\sum_{m=0}^{199}|\langle m|\psi(t_{1})\rangle|^{2}D_{m}^{(x)}=0.180$, by taking the lowest 200 lowest energy eigenstates of the equilibrium Hamiltonian, which concentrates most of the weight of $|\psi(t)\rangle$ at the time $t=t_{1}$ of the application of the probing flux — overlaps $|\langle m|\psi(t_{1})\rangle|^{2}$ are color-mapped in Fig. 1(d). This is to be contrasted against the equilibrium Drude weight, $D_{0}^{(x)}=0.026$, an order-of-magnitude smaller, furnishing evidence for photoinduced ballistic transport for this system size.
@@
Figure 3: (a) Gaussian pump field with amplitude $A_{0}=0.62$, frequency $\Omega=\omega_{26}\simeq 1.713\ t_{h}$, and temporal width $t_{d}=6.5t_{h}^{-1}$ — the same as in Fig. 1. Phase-averaged dynamics (solid curves) for a range of phases $\phi_{t}\in[0,2\pi)$ (shaded regions) are shown in (b)–(d). (b) Dynamics of the correlation functions $c_{{\bf i}_{0}}(2,0)$ and $c_{{\bf i}_{0}}(2,1)$, in respect to the reference site ${\bf i}_{0}$ marked in Fig. 1(a), exhibiting $\pi$-phase shift reversal after the pulse. (c) The zero-momentum occupation $n_{{\bf k}=0}$ exhibits a 37% enhancement in comparison to the equilibrium. (d) State projection coefficients for selected equilibrium eigenstates $|m\rangle$ (the ones showing significant weight over the dynamics), with state $|2\rangle$ gradually becoming dominant at long times, added by an enhanced overlap with state $|6\rangle$.
@@
Melting of $\pi$-phase shift.—  Having confirmed that an appropriate photoirradiation leads to not only a larger occupancy of the zero-momentum mode but also enhanced transport, we now focus on the fate of the charge $\pi$-shift mediated by the stripes. Under the same photoirradiation conditions, later detailed, and apart from the introduction of a probe vector potential, Figure 3 shows the dynamics of $n_{{\bf k}=0}$ closely — data are reported for a range of the pump’s phase shifts $\phi_{t}$, as well as their average, which aids in smoothening the short-time coherent oscillations. That the photoirradiation can be used to filter pairing in detriment of charge-stripes is apparent by the fact that while $n_{{\bf k}=0}$ dynamically increases [Fig. 3(c)], the $\pi$-phase shift disappears: $c_{{\bf i}_{0}}(2,0)$ and $c_{{\bf i}_{0}}(2,1)$ switch their signs [Fig. 3(b)] at the time at which the pump [Fig. 3(a)] has the largest amplitude.
@@
This can be understood in terms of the overlaps $|\langle m|\psi(t)\rangle|^{2}$ with low-lying equilibrium eigenstates $|m\rangle$ in Fig. 3(d). At long times, the dynamics stabilize most of the spectral weight in state $|2\rangle$, the target state, as well as $|6\rangle$ with a still relevant weight ($13\%$) with the initial state, $|0\rangle$. For that, one needs to optimize pump parameters, including the pump’s width $t_{d}$ and amplitude $A_{0}$, as indicated in Figs. 4(a) and 4(b), resulting in optimal $A_{0}=0.62$, and $t_{d}=6.5t_{h}^{-1}$. However, additional setting of the pump’s frequency is required to achieve the most accurate targeting of the excited state. Previous investigations in various models [74, 75, 76] established that a resonant photoirradiation, $\Omega=E_{\rm target}-E_{0}$, could accomplish a direct targeting of relevant excited states with eigenenergies $E_{\rm target}$. At the crux of such a process is the fact that first-order time-dependent perturbation theory stemming from the pulse application (details in the Supplemental Materials [77]) establishes that the $x$-direction current operator, $\hat{J}_{x}=it_{h}\sum_{\bf i}(\hat{b}_{\bf i+\hat{x}}^{\dagger}\hat{b}_{\bf i}^{\phantom{\dagger}}-b_{\bf i}^{\dagger}\hat{b}_{\bf i+\hat{x}}^{\phantom{\dagger}})$, connects these states with a robust $|\langle m_{\rm target}|\hat{J}_{x}|0\rangle|$. Unfortunately, our investigation shows that, in our case, the matrix elements of the current operator between $|0\rangle$ and $|2\rangle$ are small [Fig. 4(c)].
@@
Figure 4: Dependence on the pump width $t_{d}$ and amplitude $A_{0}$ of the long-time overlap at $t_{\rm max}=9t_{d}$ with selected equilibrium eigenstates, $|m=2\rangle$ (a) and $|m=6\rangle$ (b). A finer mesh resolution is shown in the area with an enhanced overlap $|\langle\psi(t_{\rm max})|2\rangle|^{2}$ while the cross marker pinpoints the maximum value. Here, the pump frequency is set at $\Omega=\omega_{26}=1.713\ t_{h}$. (c) The equilibrium $x$-direction current matrix in the low-lying spectrum; dashed lines highlight matrix elements with the target state $|2\rangle$.

### v2

@@ Abstract
The interplay between charge stripes and pairing has long been a subject of scrutiny in a broad class of unconventional superconductors, as in some cases it is unclear whether this interplay benefits the ensuing superfluidity. Experiments that explore the out-of-equilibrium dynamics of these systems aim to tip the balance toward one phase or the other by selectively coupling to relevant modes. Leveraging the fact that competition between stripes and pairing is not exclusive to fermionic systems, we explore the photoirradiation dynamics of interacting hardcore bosons, in which density-wave phase-reversal melting leads to enhanced phase-coherent transport response, as quantified by the dynamic amplification of both the zero-momentum occupancy and the condensate fraction, as well as finite out-of-equilibrium charge stiffness and superfluid weight, for a given system size. Our results, obtained using unbiased methods for an interacting system on a ladder geometry, demonstrate how one can engineer time-dependent perturbations to release suppressed orders, potentially providing insight into the underlying mechanism in related experiments.
1 Introduction
Time-dependent experiments provide another angle in the longstanding quest to understand the mechanism binding otherwise repulsive fermions, supporting the emergence of unconventional superconductivity. In particular, rapid energy insertion in the form of an ultrafast laser pump has arisen as a capable technique [1, 2] that coherently drives the dynamics of the electrons in ways that can control magnetic [3, 4], charge [5, 6, 7, 8], and superconducting orders [9, 10, 11, 12, 13, 14, 15]. Owing to the entwining of the different orders that permeate the doped cuprate physics [16, 17], such an approach is one in which coupling to certain degrees of freedom can suppress specific ordering while releasing others [2, 18].
@@ 1 Introduction
Although the physics of stripes in equilibrium, including fluctuating/incommensurate ones, is now well studied within a broad class of models and numerical methods [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], direct indication of the dynamical interplay between charge ordering and pairing is nonexistent in unbiased numerics of many-body models even though this is the precise mechanism explored in experiments [9, 43, 11, 44] – we aim to fill this void  [^1] We note microscopic models for cuprates have been studied under photoirradiation [87, 88, 89, 90], but the interplay with stripe physics is currently missing.. Our key observation is that the stripe physics and antiphase of local order are not intrinsic to repulsive interactions [46], nor to fermionic statistics [47], including the fluctuating-static dichotomy [47]. This fundamental, and highly non-trivial, analogy allows one to unbiasedly solve the dynamics of Hamiltonians that contain specific ingredients of actual cuprate physics — namely, strong interactions, superfluidity, and antiphase correlations across stripes — even if the pairing is isotropic, unlike what is observed in a large class of unconventional superconductors, and even though our model does not contain the spin degrees of freedom which are relevant for cuprate materials.
Despite these differences, our main result is that by adjusting the characteristics of the time-dependent perturbation in the form of a pump in a strongly correlated system, one can show that an effective nonlinear optical coupling is responsible for exciting the system from the ground state in a way that a static stripe with antiphase ordering melts while coherence and dissipationless-transport diagnostics are enhanced dynamically, a feature in line with experimental exploration of actual cuprates under photoirradiation.
2 Model
Our starting point is the Hubbard model for repulsive hardcore bosons,
@@ 2 Model
Figure 1: (a) Schematic illustration of Eq. (1) with relevant terms annotated, subjected to $x$-polarized photoirradiation. Here, the hardcore bosons are represented by a composite local fermion. Momentum distribution profiles $n_{\bf k}$ focusing on $k_{y}=0$ (b) and $k_{y}=\pi$ (c), comparing the equilibrium (before pump) with the out-of-equilibrium (after pump, averaged results from $t\in[11t_{d},26t_{d}]=[72,169]t_{h}^{-1}$; shading gives the standard deviation) response to the external field. We observe an average relative enhancement of about 37% for the zero-momentum condensation under the photoirradiation. (d) The $x$-direction charge stiffness resolved for the first 200 eigenstates $|m\rangle$ of $\hat{\cal H}$; marker colors are mapped by the overlap of the time evolved state at $t_{1}=5t_{d}$ and the corresponding $|m\rangle$. (e) The instantaneous energy $E(t)=\langle\psi(t)|\hat{\cal H}(t)|\psi(t)\rangle$ over the dynamics; the shaded region marks the probing for the dynamical charge stiffness starting at $t_{1}$, with a linear-in-time vector potential with slope $\delta A_{x}$ — top panel depicts the total ${\bf A}(t)$. Inset shows the energy variation in the probe regime as a function of the effective flux $\Phi_{x}(t)=\delta A_{x}L_{x}(t-t_{1})$ for $t>t_{1}$; $\delta A_{x}=0.01$. Dashed line displays an approximant at small fluxes that quantifies the dynamical charge stiffness at $t=t_{1}$, confirming a robust charge transport induced by the pump (see text). (f) The dynamics of the condensate fraction, highlighting a $\sim 34\%$ increase when considering the running average (darker curve). Pulse parameters are $\Omega=1.713\ t_{h}$, $A_{0}=0.62$ and $t_{d}=6.5t_{h}^{-1}$.
@@
3 Enhanced condensation and dissipationless transport
Interacting hardcore bosons at $T=0$ give rise to either charge ordered phases or a superfluid phase [52, 53, 54, 55, 56, 57], depending on the interaction-hopping ratio $V/t_{h}$ — in what follows we set $V/t_{h}=4$. Tendency towards phase separation and stripe formation in the vicinity of half-filling, $\langle\hat{n}\rangle=1/2$, has been reported [55, 57], and stabilization of such a regime can be made explicit via an external periodic potential [47] in the doped regime  [^4] Including an extra potential has been used in other contexts to stabilize magnetic ordering for example [91]. Notably, in this scenario, the propensity for spatial organization of holes is accompanied by a reversal in the phase of the checkerboard density modulation across a stripe boundary provided the corresponding chemical potential $V_{0}$ is sufficiently large – this is often dubbed as a $\pi$-phase shift [33, 46, 47]. Focusing on the hole density $\delta\equiv 1/2-\langle\hat{n}\rangle=1/8$, Fig. 2(a) displays the connected density-density correlations across a stripe, $c_{{\bf i}_{0}}(dx,dy)=\langle\hat{n}_{{\bf i}_{0}}\hat{n}_{{\bf i}+d{\bf r}}\rangle-\langle\hat{n}_{{\bf i}_{0}}\rangle\langle\hat{n}_{{\bf i}+d{\bf r}}\rangle$; ${\bf i}_{0}$ is a reference site and $d{\bf r}\equiv(dx,dy)$ the spatial displacement. As $V_{0}/t_{h}$ increases, the same sublattice correlation $c_{{\bf i}_{0}}(2,0)$ across a stripe transitions from positive to negative row whereas $c_{{\bf i}_{0}}(2,1)$ shows the opposite trend, establishing a direct verification of the antiphase charge ordered pattern once hole stripes are sufficiently robust.
@@ 3 Enhanced condensation and dissipationless transport
Nonetheless, condensation need not be necessarily tied to superfluid behavior, in general [60]. A necessary condition for dissipationless charge transport is given by the charge stiffness (or Drude weight), whose finite value indicates an ideal metal with ballistic transport [61, 62, 63]. For the $m$-th eigenstate of $\hat{\cal H}$ with eigenenergy $E_{m}$,
@@
The inset in Fig. 1(f) shows the time-dependence of the energy gain during the probe time. Indeed, at short times, a quadratic dependence in $\Phi_{x}(\Delta t)$ emerges, indicating a finite $D^{(x)}$ and thus dissipationless transport after the pulse. The second estimation of this physical behavior is that $D^{(x)}$ closely matches with an approximant of the out-of-equilibrium Drude weight $\tilde{D}^{(x)}=\sum_{m=0}^{199}|\langle m|\psi(t_{1})\rangle|^{2}D_{m}^{(x)}=0.180$, by taking the lowest 200 lowest energy eigenstates of the equilibrium Hamiltonian, which concentrates most of the weight of $|\psi(t)\rangle$ at the time $t=t_{1}$ of the application of the probing flux — overlaps $|\langle m|\psi(t_{1})\rangle|^{2}$ are color-mapped in Fig. 1(d). This is to be contrasted against the equilibrium Drude weight, $D_{0}^{(x)}=0.026$, an order-of-magnitude smaller, furnishing evidence for photoinduced ballistic transport for this system size.
@@
Figure 3: (a) Gaussian pump field with amplitude $A_{0}=0.62$, frequency $\Omega=\omega_{26}\simeq 1.713\ t_{h}$, and temporal width $t_{d}=6.5t_{h}^{-1}$ — the same as in Fig. 1. Phase-averaged dynamics (solid curves) for a range of phases $\phi_{t}\in[0,2\pi)$ (shaded regions) are shown in (b)–(d). (b) Dynamics of the correlation functions $c_{{\bf i}_{0}}(2,0)$ and $c_{{\bf i}_{0}}(2,1)$, in respect to the reference site ${\bf i}_{0}$ marked in Fig. 1(a), exhibiting $\pi$-phase shift reversal after the pulse — the corresponding density dynamics is shown in Appendix C. (c) The zero-momentum occupation $n_{{\bf k}=0}$ exhibits a 37% enhancement in comparison to the equilibrium. (d) State projection coefficients for selected equilibrium eigenstates $|m\rangle$ (the ones showing significant weight over the dynamics), with state $|2\rangle$ gradually becoming dominant at long times, added by an enhanced overlap with state $|6\rangle$.
4 Enhanced superfluid response
The onset of ballistic transport is a necessary, but not sufficient, condition for establishing superfluid behavior. A definitive signature requires a simultaneous enhancement of phase coherence. Figure 1(f) displays the dynamics of the condensate fraction, $f_{0}(t)=\lambda_{\rm max}(t)/N$, where $N=L_{x}\times L_{y}$ and $\lambda_{\rm max}$ is the largest eigenvalue of the single-particle density matrix $\langle\hat{b}^{\dagger}_{\bf i}\hat{b}^{\phantom{\dagger}}_{\bf j}\rangle$. In agreement with the increase in zero-momentum occupation shown in Fig. 1(b), $f_{0}(t)$ rises by approximately $34\%$ after photoirradiation, confirming that the buildup of phase coherence indeed accompanies the light-induced enhancement of condensation. This is a direct reflection of the longer-ranged correlations following the pulse — see Appendix F for an in-depth analysis, including the out-of-equilibrium estimation of the superfluid weight [Appendix G]. It is important to highlight that superfluidity entails long-range order, which can only be obtained with finite-size scalings, elusive for a calculation in out-of-equilibrium. As such, here we refer to an enhanced superfluid response on our finite-system size.
5 Melting of $\pi$-phase shift
@@
Having confirmed that an appropriate photoirradiation leads to not only a larger occupancy of the zero-momentum mode but also enhanced transport, we now focus on the fate of the charge $\pi$-shift mediated by the stripes. Under the same photoirradiation conditions, later detailed, and apart from the introduction of a probe vector potential, Figure 3 shows the dynamics of $n_{{\bf k}=0}$ closely — data are reported for a range of the pump’s phase shifts $\phi_{t}$, as well as their average, which aids in smoothening the short-time coherent oscillations. That the photoirradiation can be used to filter pairing in detriment of charge-stripes is apparent by the fact that while $n_{{\bf k}=0}$ dynamically increases [Fig. 3(c)], the $\pi$-phase shift disappears: $c_{{\bf i}_{0}}(2,0)$ and $c_{{\bf i}_{0}}(2,1)$ switch their signs [Fig. 3(b)] at the time at which the pump [Fig. 3(a)] has the largest amplitude.
@@
This can be understood in terms of the overlaps $|\langle m|\psi(t)\rangle|^{2}$ with low-lying equilibrium eigenstates $|m\rangle$ in Fig. 3(d). At long times, the dynamics stabilize most of the spectral weight in state $|2\rangle$, the target state, as well as $|6\rangle$ with a still relevant weight ($13\%$) with the initial state, $|0\rangle$. For that, one needs to optimize pump parameters, including the pump’s width $t_{d}$ and amplitude $A_{0}$, as indicated in Figs. 4(a) and 4(b), resulting in optimal $A_{0}=0.62$, and $t_{d}=6.5t_{h}^{-1}$ — Appendix D shows that these pump parameters also maximize the long-time average of $n_{{\bf k}=0}$. However, additional setting of the pump’s frequency is required to achieve the most accurate targeting of the excited state. Previous investigations in various models [73, 74, 75] established that a resonant photoirradiation, $\Omega=E_{\rm target}-E_{0}$, could accomplish a direct targeting of relevant excited states with eigenenergies $E_{\rm target}$. At the crux of such a process is the fact that first-order time-dependent perturbation theory stemming from the pulse application (details in Appendix A) establishes that the $x$-direction current operator, $\hat{J}_{x}=it_{h}\sum_{\bf i}(\hat{b}_{\bf i+\hat{x}}^{\dagger}\hat{b}_{\bf i}^{\phantom{\dagger}}-b_{\bf i}^{\dagger}\hat{b}_{\bf i+\hat{x}}^{\phantom{\dagger}})$, connects these states with a robust $|\langle m_{\rm target}|\hat{J}_{x}|0\rangle|$. Unfortunately, our investigation shows that, in our case, the matrix elements of the current operator between $|0\rangle$ and $|2\rangle$ are small [Fig. 4(c)].
@@
Figure 4: Dependence on the pump width $t_{d}$ and amplitude $A_{0}$ of the long-time overlap at $t_{\rm max}=9t_{d}$ with selected equilibrium eigenstates, $|m=2\rangle$ (a) and $|m=6\rangle$ (b). A finer mesh resolution is shown in the area with an enhanced overlap $|\langle\psi(t_{\rm max})|2\rangle|^{2}$ while the cross marker pinpoints the maximum value. Here, the pump frequency is set at $\Omega=\omega_{26}=1.713\ t_{h}$. (c) The equilibrium $x$-direction current matrix in the low-lying spectrum; dashed lines highlight matrix elements with the target state $|2\rangle$. Other relevant matrix elements for the perturbative analysis are shown in Appendix B.
6 The non-linear optical coupling